# A mesh-free hybrid Chebyshev-Tucker tensor format with applications to multi-particle modelling

### v1

@@ 3.2.2 Algebraic case
The above result in fact holds for any dimension $d\geq 3$. Therefore, we will prove the lemma in the general $d$-dimensional case as follows.
Proof.
Recall that by construction, $q_{k}^{(\ell)}$ interpolates the data ${\bf a}_{k}^{(\ell)}$ on the uniform grid $\{t_{i_{\ell}}^{(\ell)}\}_{i_{\ell}=1}^{n_{\ell}}$ and the Chebyshev interpolants have uniform error $\delta_{k}^{(\ell)}$. Therefore, for each $\ell=1,\cdots,d$, $k=1,\cdots,R$ and $i_{\ell}=1,\cdots,n_{\ell}$, we have
$|{\bf a}_{k}^{(\ell)}(i_{\ell})-g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})|=|q_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})-g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})|\leq\|q_{k}^{(\ell)}-g_{k}^{(\ell)}\|_{\infty}=\delta_{k}^{(\ell)}$ (3.23)
and in particular,
$|g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})|\leq|{\bf a}_{k}^{(\ell)}(i_{\ell})|+\delta_{k}^{(\ell)}\leq\|{\bf a}_{k}^{(\ell)}\|_{\infty}+\delta_{k}^{(\ell)}\leq 1+\delta_{k}^{(\ell)},$ (3.24)
where we have used the fact that the canonical vectors are $\ell_{2}$ normalized. Therefore, for any $i_{1},\cdots,i_{d}$, we have
$\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},\cdots,t_{i_{d}}^{(d)})-{\bf F}(i_{1},\cdots,i_{d})\right|$ (3.25)
$=$ $\left|\sum_{k=1}^{R}\xi_{k}g_{k}^{(1)}(t_{i_{1}}^{(1)})\cdots g_{k}^{(d)}(t_{i_{d}}^{(d)})-\sum_{k=1}^{R_{L}}\xi_{k}{\bf a}_{k}^{(1)}(i_{1})\cdots{\bf a}_{k}^{(d)}(i_{d})\right|$
$\leq$ $\sum_{k=1}^{R_{L}}\xi_{k}\left|g_{k}^{(1)}(t_{i_{1}}^{(1)})\cdots g_{k}^{(d)}(t_{i_{d}}^{(d)})-{\bf a}_{k}^{(1)}(i_{1})\cdots{\bf a}_{k}^{(d)}(i_{d})\right|$
$=$ $\sum_{k=1}^{R_{L}}\xi_{k}\left|\sum_{\ell=1}^{d}\left(\prod_{r=1}^{\ell-1}g_{k}^{(r)}(t_{i_{\ell}}^{(r)})\right)\left(g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})-{\bf a}_{k}^{(\ell)}(i_{\ell})\right)\left(\prod_{r=\ell+1}^{d}{\bf a}_{k}^{(r)}(i_{r})\right)\right|$
$\leq$ $\sum_{k=1}^{R}\xi_{k}\sum_{\ell=1}^{d}\prod_{r=1}^{\ell-1}\left(1+\delta_{k}^{(r)}\right)\delta_{k}^{(\ell)}.$
This completes the proof of the first inequality. The second inequality follows from
$\sum_{k=1}^{R}\xi_{k}\sum_{\ell=1}^{d}\prod_{r=1}^{\ell-1}\left(1+\delta_{k}^{(r)}\right)\delta_{k}^{(\ell)}\leq\sum_{k=1}^{R_{L}}\xi_{k}\sum_{\ell=1}^{d}\left(1+\delta\right)^{\ell-1}\delta=\|\bm{\xi}\|_{1}\left((1+\delta)^{d}-1\right).$ (3.26)
The bound Eq. 3.22 shows that the error of the functional canonical tensor format is controlled by the $\ell_{\infty}$ error of the univariate Chebyshev interpolants $\delta$ and goes to zero as $\delta\to 0$ but is amplified exponentially. However, by assuming $\delta<1/d$, which can be satisfied for functions with reasonable regularity, the error of the functional canonical tensor format only grows linearly with the dimension $d$.
@@
If $\delta<1$, we have
$\max_{i_{1},i_{2},i_{3}}\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\|\bm{\xi}\|_{1}\left(\delta d(1-\delta)+\delta^{2}(\mathrm{e}^{d}-1)\right).$ (3.27)
If $\delta d<1$, we have
$\max_{i_{1},i_{2},i_{3}}\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\|\bm{\xi}\|_{1}\min\left\{\dfrac{\delta d}{1-\delta d},(\mathrm{e}-1)\delta d\right\}.$ (3.28)
Proof.
For $\delta<1$, we have
$(1+\delta)^{d}-1\leq\mathrm{e}^{\delta d}-1.$ (3.29)
Then the second inequality follows from the following elementary inequalities for the exponential function for $x\in(0,1)$:
$\mathrm{e}^{x}\leq(\mathrm{e}-1)x+1\text{ and }\mathrm{e}^{x}\leq\dfrac{1}{1-x}.$ (3.30)
For the first inequality, let $f(\delta,d)=\delta d(1-\delta)+\delta^{2}(\mathrm{e}^{d}-1)-\mathrm{e}^{\delta d}+1$, then $f\geq 0$ for $10<\delta<$ 1 and $d>0$ follows from standard calculus arguments. ∎
Now, we consider the error introduced by the RHOSVD truncation [31] in Algorithm 4, namely, the second term in the right-hand side of Eq. 3.19.
Lemma 3.8.
With the above notations, we have
$\|\hat{f}_{\bf m}-\tilde{f}_{\bf m}\|_{\infty}\leq\sqrt{m_{1}m_{2}m_{3}}\|\bm{\xi}\|\sum_{\ell=1}^{3}\left(\sum_{k=r_{\ell}+1}^{\min\{m_{\ell},R\}}(\sigma_{k}^{(\ell)})^{2}\right)^{1/2},$ (3.31)
where $\|\bm{\xi}\|^{2}=\sum_{k=1}^{R}\xi_{k}^{2}$ is the $\ell_{2}$ norm of the normalization constants.
Proof.
By the same arguments as in the proof of Proposition 3.4, we have
$\|\tilde{f}_{\bf m}-\hat{f}_{\bf m}\|_{\infty}\leq\sqrt{m_{1}m_{2}m_{3}}\|{\bf C}-\hat{\bf{C}}\|_{\mathrm{F}}.$ (3.32)
The desired result then follows from the standard error analysis of the RHOSVD truncation [31, Theorem 2.5]. See also [27] for a more recent presentation. ∎
Combining Lemmas 3.5 and 3.8 and the triangle inequality Eq. 3.19, we arrive at the following theorem immediately, which provides a bound for the error of the ChebTuck approximation in the algebraic case of CP tensor input.
Theorem 3.9.
$\max_{i_{1},i_{2},i_{3}}\left|\hat{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq$ $\|\bm{\xi}\|_{1}\left((1+\delta)^{3}-1\right)$ (3.33)
$+$ $\sqrt{m_{1}m_{2}m_{3}}\|\bm{\xi}\|\sum_{\ell=1}^{3}\left(\sum_{k=r_{\ell}+1}^{\min\{m_{\ell},R\}}(\sigma_{k}^{(\ell)})^{2}\right)^{1/2}.$
@@
From the above theorem, one can design an adaptive algorithm that constructs the ChebTuck approximation of a CP tensor with a given error $\varepsilon>0$ by choosing the smallest possible $r_{\ell}$ in each mode $\ell=1,2,3$ such that the error bound is less than $\varepsilon$. This is because all the terms in the bound are computable, for example, $\delta$ will be anyway computed when we construct the univariate Chebyshev interpolants, and the singular values $\sigma_{k}^{(\ell)}$ are available after the SVD decomposition of the side matrices in the CP format.
3.3 Complexity bounds for the Algorithms 1 - 4
In this section, we briefly discuss the complexity of the proposed algorithms for the ChebTuck approximation of a function in both the functional and algebraic cases.
@@ 3.3 Complexity bounds for the Algorithms 1 - 4
For the functional case Algorithm 1, let $C_{f}$ be the cost of one functional call. For the algebraic cases Algorithms 2, 3 and 4, let $n=\max\{n_{1},n_{2},n_{3}\}$ be the maximum grid size. For the algebraic cases with CP tensor input Algorithms 3 and 4, let $R$ be the rank of the CP tensor. Let $m=\max\{m_{1},m_{2},m_{3}\}$ be the maximum Chebyshev degree and $\varepsilon>0$ be the Tucker-ALS or RHOSVD tolerance. Then the complexity of the algorithms is estimated by
@@
Algorithm 2: $C_{4}m^{3}(m+n)+C_{5}m^{3}+C_{1}m^{3}\log m+C_{2}m^{4}+C_{3}m^{4}|\log\varepsilon|$.
Algorithm 3: $C_{6}nR+C_{7}mR+C_{8}m^{3}R+C_{2}m^{4}+C_{3}m^{4}|\log\varepsilon|$.
Algorithm 4: $C_{6}nR+C_{7}mR+C_{9}mR\log m+C_{10}mR\min\{m,R\}+C_{11}R|\log\varepsilon|^{3}$.
Proof.
@@
For the functional case Algorithm 1, the number of functional calls is $m^{3}$ and therefore the cost of 3 is $C_{f}m^{3}$. The cost of 4 is $C_{1}m^{3}\log m$ by the FFT-based Chebyshev coefficient computation. The cost of Tucker-ALS is $C_{2}m^{4}+C_{3}m^{4}r$ by the standard results on its complexity, see e.g. [30, Chapter 3.4.3], where the Tucker rank $r=\mathcal{O}(|\log\varepsilon|)$ is assumed. This assumption is true for a wide range of functions including Newton kernel in our numerical experiments [28]. This completes the proof for the functional case.
@@
For the algebraic cases, we first claim that the cost of evaluating a cubic spline interpolation at $m$ points given data on uniform $n$-point grid is $\mathcal{O}(m+n)$. Indeed, the computation of the spline coefficients requires solving a tri-diagonal system of size $44n-$ 4, which costs $\mathcal{O}(n)$, and the evaluation of the spline at $m$ points costs $\mathcal{O}(m)$ in case the $n$-point grid is uniform. Similarly, the cost of trivariate cubic spline interpolation which is required in Algorithm 2 is $\mathcal{O}(m^{3}(m+n))$. This then completes the proof for the complexity of Algorithm 2.
@@
For the complexity of Algorithm 3, the computation of the function evaluation tensor ${\bf T}$ requires $\mathcal{O}(m^{3}R)$ operations, explaining the third term in the estimate. Finally, for the complexity of Algorithm 4, the computation of the canonical vectors ${\bf a}_{k}^{(\ell)}$ requires $\mathcal{O}(m\log m)$ operations by using FFT. The complexity of the RHOSVD truncation is the sum of the costs of the truncated SVD of the side matrices, amounting to $\mathcal{O}(mR\min\{m,R\})$, and the cost of reconstruction of the Tucker core tensor $\bm{\beta}$, which is $\mathcal{O}(Rr^{3})$, see [27] for more details. ∎
We remark that the main benefits of the proposed tensor techniques are due to the following features.
  • A
    The mesh-free tensor approximation of a function in $\mathbb{R}^{d}$ that requires evaluation of the function of interest only at a minimal number of interpolation points.
  • B
    Calculation and storage of the discrete approximation for the multivariate function in $\mathbb{R}^{d}$ over huge $n\times n\times n$ grid is not required, as well as subsequent numerical Tucker-ALS approximation of the resultant 3D tensor in $\mathbb{R}^{n\times n\times n}$ having rather complicated landscape.
  • C
    Cheap and fast Tucker-ALS rank optimization for the small $m\times m\times m$ coefficient tensor with $m\ll n$.
  • D
    In case of functional hybrid approximation one obtains the low-rank Tucker/CP tensors that approximate the initial function by discretizing the product Chebyshev polynomials on a given fine $n^{\otimes 3}$ tensor grid, thus avoiding the storage and tensor decomposition of the tensor in $\mathbb{R}^{n\times n\times n}$, serving for the full grid-based representation of a function $f(x)$.
We recall that our main goals in this paper is to construct the method that allows to compute the mesh-less Tucker tensor approximation of the target function with the almost optimal tensor rank ${\bf r}$ which is usually much smaller than the Chebyshev polynomials degree $m$. The computed ChebTuck format can be used for further multilinear algebra calculations within the considered computational tasks, in particular, in tensor numerical methods for solving various PDEs arising in large-scale scientific computing. The Tucker rank of the output tensor is the most important parameter for the complexity characteristics of the subsequent multilinear tensor calculus in the respective applications.
@@
However, we notice that there is the traditional trade-off between the accuracy and computational complexity of the numerical algorithms. In this concern, we comment that given $\varepsilon>0$, the polynomial degree $m=m(\varepsilon)$ depends on the regularity of input function $f(x)$ and may increase dramatically for non-regular functions having, for example, multiple local singularities or cusps.
This is the case for interaction potentials of many-particle systems typically arising in quantum computational chemistry and bio-molecular modeling. For such challenging problems, the successful application of our method is based on the use of range-separated (RS) tensor formats. The efficiency of such an approach will be demonstrated in what follows.
4 Applications to multi-particle modelling

### v2

@@ 3.2.2 Algebraic case
Proof 3.5.
Recall that by construction, $q_{k}^{(\ell)}$ interpolates the data ${\bf a}_{k}^{(\ell)}$ on the uniform grid $\{t_{i_{\ell}}^{(\ell)}\}_{i_{\ell}=1}^{n_{\ell}}$ and the Chebyshev interpolants have uniform error $\delta_{k}^{(\ell)}$. Therefore, for each $\ell=1,\ldots,d$, $k=1,\ldots,R$ and $i_{\ell}=1,\ldots,n_{\ell}$, we have
(25) $|{\bf a}_{k}^{(\ell)}(i_{\ell})-g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})|=|q_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})-g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})|\leq\|q_{k}^{(\ell)}-g_{k}^{(\ell)}\|_{\infty}=\delta_{k}^{(\ell)}$
and in particular,
(26) $|g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})|\leq|{\bf a}_{k}^{(\ell)}(i_{\ell})|+\delta_{k}^{(\ell)}\leq\|{\bf a}_{k}^{(\ell)}\|_{\infty}+\delta_{k}^{(\ell)}\leq 1+\delta_{k}^{(\ell)},$
where we have used the fact that the canonical vectors are $\ell_{2}$ normalized. Therefore, for any $i_{1},\ldots,i_{d}$, by the definition of $\tilde{f}_{\bf m}$ in Eq. 23 and $\bf F$} in Eq. 18, we have
$\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},\ldots,t_{i_{d}}^{(d)})-{\bf F}_{i_{1},\ldots,i_{d}}\right|\leq{\sum}_{k=1}^{R}\xi_{k}\left|\prod_{\ell=1}^{d}g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})-\prod_{\ell=1}^{d}{\bf a}_{k}^{(\ell)}(i_{\ell})\right|$
$\leq$ $\sum_{k=1}^{R}\xi_{k}\sum_{\ell=1}^{d}\prod_{r=1}^{\ell-1}\left(1+\delta_{k}^{(r)}\right)\delta_{k}^{(\ell)}\leq\sum_{k=1}^{R}\xi_{k}\sum_{\ell=1}^{d}\left(1+\delta\right)^{\ell-1}\delta=\|\bm{\xi}\|_{1}\left((1+\delta)^{d}-1\right),$
where the second inequality results from applying a standard telescopic argument.
By assuming $\delta<1/d$, which is easily satisfied in practical computations, we can further improve the bound Eq. 24 to a linear growth with respect to the dimension $d$ as in the following corollary.
Corollary 3.6.
Under the assumptions of Lemma 3.4, further assuming $\delta d<1$ yields
(27) $\max_{i_{1},i_{2},i_{3}}\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\|\bm{\xi}\|_{1}\min\left\{\dfrac{\delta d}{1-\delta d},(\mathrm{e}-1)\delta d\right\}.$
Proof 3.7.
First note that $(1+\delta)^{d}\leq\mathrm{e}^{\delta d}$. Then the desired result follows from the elementary inequalities $\mathrm{e}^{x}\leq(\mathrm{e}-1)x+1\text{ and }\mathrm{e}^{x}\leq(1-x)^{-1}$ for $x\in(0,1)$.
The following lemma states that the error of the univariate Chebyshev interpolation of the cubic spline $q_{k}^{(\ell)}$ decays cubically with respect to the Chebyshev degree $m$.
Lemma 3.8.
Let $V_{k}^{\ell}<+\infty$ be the bounded variation of the third derivative of the spline $q_{k}^{(\ell)}$ and $V=\max_{k,\ell}V_{k}^{\ell}$. Then we have
(28) $\delta_{k}^{(\ell)}\leq\dfrac{4V_{k}^{\ell}}{3\pi(m-3)^{3}}\text{ for }k=1,\ldots,R,\quad\ell=1,2,3\text{ and }\delta\leq\dfrac{4V}{3\pi(m-3)^{3}}.$
Proof 3.9.
By definition, $q_{k}^{(\ell)}$ is a $C^{2}$ piecewise cubic polynomial, and thus absolutely continuous differentiable until the second derivative. The third derivative is piecewise constant thus of bounded variation $V_{k}^{\ell}$. Note that $V_{k}^{\ell}$ is uniquely determined by the spline knots (here the uniform grid $\{t_{i_{\ell}}^{(\ell)}\}_{i_{\ell}=1}^{n_{\ell}}$) and the spline values ${\bf a}_{k}^{(\ell)}(i_{\ell})$. The results then follow from the standard error bound for Chebyshev interpolation of differentiable functions, see e.g., [59, Thm. 7.2].
Now, we consider the error introduced by the RHOSVD truncation [39] in Algorithm 3, namely, the second term in the right-hand side of Eq. 22.
Lemma 3.10.
Let $\sigma_{k}^{(\ell)}$, $k=1,\ldots,\min\{m,R\}$ be the descending singular values of the side matrices $C^{(\ell)}=[{\bf c}_{1}^{(\ell)},\cdots,{\bf c}_{R}^{(\ell)}]\in\mathbb{R}^{m\times R}$ of the Tucker format Eq. 20, and $r_{\ell}$ be the truncation rank of the RHOSVD approximation, then we have
(29) $\|\hat{f}_{\bf m}-\tilde{f}_{\bf m}\|_{\infty}\leq m^{d/2}\|\bm{\xi}\|\sum_{\ell=1}^{d}\Big(\sum_{k=r_{\ell}+1}^{\min\{m,R\}}(\sigma_{k}^{(\ell)})^{2}\Big)^{1/2},\text{ where }\|\bm{\xi}\|^{2}=\sum_{k=1}^{R}\xi_{k}^{2}.$
Proof 3.11.
By similar arguments as in the proof of Proposition 3.2, we have $\|\tilde{f}_{\bf m}-\hat{f}_{\bf m}\|_{\infty}\leq m^{d/2}\|{\bf C}-\hat{\bf{C}}\|_{\mathrm{F}}$. The result then follows from the standard error analysis of the RHOSVD truncation [39, Thm. 2.5]. See also [35] for a more recent presentation.
Combining Corollary 3.6, Lemma 3.10 and the triangle inequality Eq. 22, we arrive at the following theorem immediately, which provides a bound for the error of the ChebTuck approximation in the algebraic case of CP tensor input.
Theorem 3.12.
With the notations above, the error of the ChebTuck approximation $\hat{f}_{\bf m}$ of a CP tensor $\bf F$} produced by Algorithm 3 is bounded by
$\max_{i_{1},i_{2},i_{3}}\left|\hat{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\|\bm{\xi}\|_{1}\mathrm{e}\delta d+m^{d/2}\|\bm{\xi}\|\sum_{\ell=1}^{d}\Big(\sum_{k=r_{\ell}+1}^{\min\{m,R\}}(\sigma_{k}^{(\ell)})^{2}\Big)^{1/2}.$
@@
From the above theorem, one can design an adaptive algorithm that constructs the ChebTuck approximation of a CP tensor with a given error $\varepsilon>0$ by choosing the smallest possible $r_{\ell}$ in each mode $\ell=1,2,3$ such that the error bound is less than $\varepsilon$. This is because all the terms in the bound are computable. Specifically, $\delta$ will be anyway computed when we construct the univariate Chebyshev interpolants, and the singular values $\sigma_{k}^{(\ell)}$ are available after the SVD decomposition of the side matrices in the CP format.
3.2.3 The case of multi-particle potentials
In this section, we consider the ChebTuck approximation error for multi-particle potentials of the form Eq. 33, which is also the main application area in our numerical experiments. We show that with the use of the range-separated tensor format [6], the error of the ChebTuck approximation of the long-range part decays exponentially fast with respect to the truncation rank $r_{\ell}$, i.e., we have the following Corollary 3.14 for the second term of the error bound in Theorem 3.12.
@@ 3.2.3 The case of multi-particle potentials
Suppose ${\bf F}\in\mathbb{R}^{n\times n\times n}$ is the tensor representation of the long-range part of an $N$-particle potential satisfying the conditions of [6, Thm. 3.1]. Then the singular values of the side matrices $C^{(\ell)}$ decay exponentially fast, resulting in the following error bound
(30) $\sum_{\ell=1}^{d}\Big(\sum_{k=r+1}^{\min\{m,R\}}(\sigma_{k}^{(\ell)})^{2}\Big)^{1/2}\leq C_{m,n}Nd\exp\left(-\exp\left(r^{\frac{2}{3}}\right)\right),$
where, given the target function $f$, the constant $C=C_{m,n}$ depends only on the Chebyshev degree $m$ and the size $n$ of the uniform grid. In particular, the $\varepsilon$-Tucker rank of the canonical CCT tensor $\bf C$} is estimated by $\mathcal{O}(\log^{3/2}|\log(\varepsilon/C_{m,n}N)|)$.
Proof 3.15.
Let $A^{(\ell)}=[{\bf a}_{1}^{(\ell)},\cdots,{\bf a}_{R}^{(\ell)}]\in\mathbb{R}^{n\times R}$ be the side matrix of the CP tensor $\bf F$} Eq. 18. Then [6, Thm. 3.1] implies that the $\varepsilon$-rank of the side matrix $A^{(\ell)}$ has the bound $r_{\varepsilon}(A^{\ell})\leq C\log^{3/2}(|\log(\varepsilon/N)|)$ thus we have
(31) $\Big(\sum_{k=r_{\ell}+1}^{\min\{n,R\}}(\sigma_{k}^{(\ell)}(A^{(\ell)}))^{2}\Big)^{1/2}\leq CN\exp\left(-\exp\left(r^{\frac{2}{3}}\right)\right).$
Let $P^{(\ell)}:\mathbb{R}^{n}\to\mathbb{R}^{m}$, ${\bf q}_{k}^{(\ell)}=P^{(\ell)}{\bf a}_{k}^{(\ell)}$ be the operator that maps the canonical vector ${\bf a}_{k}^{(\ell)}$ to the evaluations of the spline $q_{k}^{(\ell)}$ at the Chebyshev points, i.e., the procedure in line 3 of Algorithm 3. It can be easily verified that the $P^{(\ell)}$ is linear thus can be represented as a matrix in $\mathbb{R}^{m\times n}$. Therefore, we have ${\bf c}_{k}^{(\ell)}=W^{(\ell)}P^{(\ell)}{\bf a}_{k}^{(\ell)}$ by line 4 of Algorithm 3, hence ${C}^{(\ell)}=W^{(\ell)}P^{(\ell)}A^{(\ell)}$. Therefore we have $\sigma_{k}^{(\ell)}\leq\|W^{(\ell)}P^{(\ell)}\|\sigma_{k}^{(\ell)}(A^{(\ell)})$ for $k=1,\ldots,R$ and $\ell=1,2,3$ by the min-max characterization of singular values. It is clear that $\|W^{(\ell)}P^{(\ell)}\|$ only depends on $m$ and $n$ since the spline knots and the Chebyshev points are only defined by $m$ and $n$ in the fixed interval [1,1][-1,1]. Absorbing the constant $C$ and the factor $\|W^{(\ell)}P^{(\ell)}\|$ into $C_{m,n}$ completes the proof.
@@
Regarding the constant $C_{m,n}$, numerical estimates show that it is in fact decaying polynomially with respect to both $m$ and $n$, though we do not have a proof for this claim. Combining Corollary 3.14 with Theorem 3.12 and Lemma 3.8, we can conclude that for the long-range part of the $N$-particle potential, the ChebTuck approximation error has the following estimate
(32) $\max_{i_{1},i_{2},i_{3}}\left|\hat{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\lesssim dm^{-3}+m^{d/2}Nd\exp\left(-\exp\left(r^{\frac{2}{3}}\right)\right).$
For fixed dimension $d$, in particular for practical applications to multi-particle potentials where $d=3$, the doubly exponential decay rate in the truncation rank $r$ essentially compensates the growth $m^{d/2}$ in the second term, resulting in a very favorable error bound. The results are consistent with our numerical experiments in Section 4.2.
3.3 Complexity bounds for Algorithms 1, 2, and 3
In this section, we briefly discuss the time complexity of the proposed algorithms for the ChebTuck approximation of a function in both functional and algebraic cases.
@@ 3.3 Complexity bounds for Algorithms 1, 2, and 3
For the functional case Algorithm 1, let $C_{f}$ be the cost of one functional call. For the algebraic cases Algorithms 2 and 3, let $n=\max_{\ell}n_{\ell}$ be the maximum grid size. For the algebraic cases with CP tensor input Algorithms 3 and 3, let $R$ be the rank of the CP tensor. Let $m$ be Chebyshev degree in each mode and $\varepsilon>0$ be the Tucker-ALS or RHOSVD tolerance. Then the complexity of the algorithms is estimated by
@@
Algorithm 2: $C_{4}m^{3}+C_{5}n^{3}+C_{1}m^{3}\log m+C_{2}m^{4}+C_{3}m^{4}|\log\varepsilon|$.
Algorithm 3: $C_{5}nR+C_{6}mR+C_{7}mR\log m+C_{8}mR\min\{m,R\}+C_{9}R|\log\varepsilon|^{3}$.
Proof 3.18.
@@
For the functional case Algorithm 1, the number of functional calls is $m^{3}$ and therefore the cost of evaluating $\bf T$} is $C_{f}m^{3}$. The cost of computing the CCT $\bf C$} is $C_{1}m^{3}\log m$ by the FFT-based Chebyshev coefficient computation. The cost of Tucker-ALS is $C_{2}m^{4}+C_{3}m^{4}r$ by the standard results on its complexity, see e.g. [37, Chapter 3.4.3], where the Tucker rank $r=\mathcal{O}(|\log\varepsilon|)$ is assumed. This assumption is true for a wide range of functions including Newton kernel in our numerical experiments [38]. This completes the proof for the functional case.
@@
For the algebraic cases, we first claim that the cost of evaluating a cubic spline interpolation at $m$ points given data on uniform $n$-point grid is $\mathcal{O}(m+n)$. Indeed, the computation of the spline coefficients requires solving a tridiagonal system of size 4n44n-4, which costs $\mathcal{O}(n)$, and an evaluation of the spline at $m$ points costs $\mathcal{O}(m)$ in case the $n$-point grid is uniform. Similarly, the cost of trivariate cubic spline interpolation which is required in Algorithm 2 is $\mathcal{O}(m^{3}+n^{3})$. This then completes the proof for the complexity of Algorithm 2.
@@
Finally, for the complexity of Algorithm 3, the first two terms correspond to the cost of line 3. The computation of the canonical vectors ${\bf c}_{k}^{(\ell)}$ in line 4 requires $\mathcal{O}(m\log m)$ operations by using FFT, explaining the third term in the estimate. The complexity of the RHOSVD truncation is the sum of the costs of the truncated SVD of the side matrices, amounting to $\mathcal{O}(mR\min\{m,R\})$, and the cost of reconstruction of the Tucker core tensor $\bm{\beta}$, which is $\mathcal{O}(Rr^{3})$, see [35] for more details.
@@
We notice that there is the traditional trade-off between the accuracy and computational complexity of the numerical algorithms. With regard to this concern, we comment that given $\varepsilon>0$, the polynomial degree $m=m(\varepsilon)$ depends on the regularity of input function $f(x)$ and may increase dramatically for non-regular functions having, for example, multiple local singularities or cusps, see for example Table 1. This is the case for interaction potentials of many-particle systems typically arising in quantum computational chemistry and bio-molecular modeling. For such challenging problems, the successful application of our method is based on the use of range-separated (RS) tensor formats, which will be discussed in what follows.
4 Applications to multi-particle modeling